# Xenon-metal pair formation in UO2 investigated using DFT+$U$

### v1

@@ Abstract
A recent experimental study on a spent uranium dioxide (UO2) fuel sample from Belgium Reactor 3 (BR3) identified a unique pair structure formed by the noble metal phase (NMP) and fission gas (xenon [Xe]) precipitate. However, the fundamental mechanism behind this structure remains unclear. The present study aims to provide a comprehensive understanding of the interaction between five different metal precipitates (molybdenum [Mo], ruthenium [Ru], palladium [Pd], technetium [Tc], and rhodium [Rh]) and the Xe fission gas atoms in UO2, by using density functional theory (DFT) in combination with the Hubbard $U$ correction to compute the formation energies involved. All DFT+$U$ calculations were performed with occupation matrix control to ensure antiferromagnetic ordering of UO2. The calculated formation energies of the Xe and solid fission products in the NMP reveal that these metal precipitates form stable structures with Xe in the following order: Mo $>$ Tc $>$ Ru $>$ Pd $>$ Rh. Notably, the formation energy of Xe-metal pairs is lower than that of the isolated single defects in all instances, with Mo showing the most negative formation energy, likely accounting for the observed pair structure formation.
@@ I Introduction
Recently, we performed advanced characterization on the spent UO2 fuel sample, whose average burnup was approximately 40 MWdkg-1, from Belgium Reactor 3 (BR3) [27, 28] and discovered that the fission product Xe forms intriguing pair structures with NMP. However, the fundamental mechanism behind these intriguing pair structures remains unknown. Although several studies have explored the stability of the NMP, the interactions between NMP and fission gases remain poorly understood. To address this knowledge gap, we employed first-principles calculations to investigate possible defect pair formations in UO2. Theoretical studies have previously utilized DFT simulations to understand defect energetics and diffusion in actinide dioxides [29, 30, 31, 32, 33]. For instance, Shilpa et al. [29] recently examined the formation and migration of various neutral defects such as vacancies, interstitials, antisites, Schotkky defects, and Frenkel defects in actinide oxides. Thompson et al. [33] conducted DFT+$U$ calculations to explore the energetics of various defects in UO2, including noble gases (He, Ne, Ar, Kr, Xe), Schottky defects, and the interaction between these defects. Andersson et al. [30] proposed mechanisms governing fission gas evolution by analyzing Xe solution thermodynamics, migration barriers, and the interaction of dissolved Xe atoms with uranium, demonstrating that Xe diffusion predominantly occurs via a vacancy-mediated mechanism. Hong et al. [32] used DFT to investigate the solubility and clustering of Ru fission products in UO2 and suggested that the probable nucleus of metallic precipitates in UO2 is metallic dimers in Schottky defects. In line with work by Hong et al. [32], we utilized the DFT+$U$ approach to understand the formation of various metal/fission-gas pair structures. Specifically, we characterized the formation energies and charge transitions of each Xe-metal pair structure, including metals such as Tc, Mo, Pd, Ru, and Rh. For the 5MPs, our DFT+$U$ calculations identified the elements that are most energetically favorable to forming these pair structures. We anticipate that our results will inform future discussions regarding the impact of defect formation on the performance and stability of UO2.
@@ III Simulation details
The chemical potential $\mu_{i}$ includes the zero-point vibrational energy and the temperature and pressure dependence of the chemical potential. By neglecting the zero-point vibration and the pressure effects, the term $\mu_{i}$ can be ignored. This approximation is reasonable, particularly for defect elements at zero temperature [52, 53, 54, 32, 55]. Thus, for defect elements, the third term in Equation (1) is just the DFT-calculated energy for each species. Furthermore, at equilibrium, the chemical potentials of U and O are related by the formation enthalpy of bulk UO2:
@@ IV Results and Discussion
To understand the formation energy of Xe-metal pairs, we first calculated the formation energy of single defects. The single defects in UO2 were constructed by replacing one U atom with either a Xe, Mo, Ru, Rh, Pd, or Tc atom. The relaxed-structure schematics for the single-atom Xe and Mo substitutions are shown in Fig. 2 (a), with all other single-defect relaxed structures being provided in the SI. The formation energy in an O-rich environment, as calculated using Equation 1 for all single defects considered in this work, is shown in Fig. 4. Among all the single defects analyzed, the Mo atom exhibits the lowest formation energy, indicating that the smallest energy penalty is incurred when substituting Mo at a U site. This is followed by Tc, Ru, Rh, Pd, and Xe. Following the analysis of single defects, we calculated the formation energies of dimer defects of the same species, with two atoms of Xe, Mo, Ru, Rh, Pd, or Tc being substituted at two neighboring U sites. The calculated formation energies of the dimers are listed in Fig. 4 (b) (denoted as “coupled”), in comparison with the formation energies measured when the two defects are spaced infinitely far apart (“isolated”). The total formation energy of the two isolated defects was estimated by summing together the formation energies of each individual defect.
@@
Figure 4 (b) clearly demonstrates that coupled dimers are energetically more favorable than dispersed atoms. Specifically, the XeU-XeU, RuU-RuU, PdU-PdU, and TaU-TaU pairs show a reduction in formation energy when in close proximity. In contrast, the changes in formation energy for RhU-RhU and MoU-MoU are marginal. These results indicate that metal atoms tend to form metallic clusters in UO2 fuel, rather than existing as dispersed atoms. Finally, we computed the formation energies of defect pairs involving Xe and various metals (Mo, Ru, Rh, Tc, and Pd) and compared these with the formation energies pertaining to when the metal and Xe atoms are isolated, as illustrated in Fig. 4 (c). Our results demonstrate that the formation energies of Xe-metal pairs are consistently lower than those of isolated defects, suggesting Xe-metal pair defects to be energetically favorable. Among the metals studied, the Xe-Mo pair exhibited the lowest formation energy, followed sequentially by Tc, Ru, Pd, and Rh. This finding indicates that, of the metals investigated in this study, Mo is the most likely candidate to form a stable pair structure with Xe.
@@
Figure 4: (a) Calculated formation energies of single defects at the U site, in the following order: Xe, Mo, Ru, Rh, Pd, and Tc. (b) Pair formation energies of dimers of the same species (blue), as compared to when these defects are isolated (orange). (c) Pair formation energies of dimers of the Xe-metal (blue), as compared to when these defects are isolated (orange).
@@ V Conclusion
In conclusion, we employed DFT+$U$ calculations to unravel the fundamental mechanism underlying the formation of Xe-M pair structures in spent UO2 fuel samples from BR3. Our study calculated the formation energies of five Xe-metal pairs, both in close proximity and in isolation. The results revealed that the formation energies of Xe-metal pairs are consistently lower than those of single defects, indicating enhanced stability for paired configurations. The stability of the Xe-M pairs follows the order of Mo $>$ Tc $>$ Ru $>$ Pd $>$ Rh, clearly highlighting that Mo is the most favorable metal for forming stable structures with Xe among the five metals considered. This elucidation of the pair formation mechanism significantly advances our understanding of fission product behavior in nuclear fuels. Future energy-dispersive X-ray spectroscopy analyses of the BR3 samples are planned to further substantiate these findings through experimental validation. These insights enhance our comprehension of defect interactions in nuclear fuel matrices.

### v2

@@ Abstract
A recent experimental study on a spent uranium dioxide (UO2) fuel sample from Belgium Reactor 3 (BR3) identified a unique pair structure formed by the noble metal phase (NMP) and fission gas (xenon [Xe]) precipitate. However, the fundamental mechanism behind this structure remains unclear. The present study aims to provide an understanding of the interaction between five different metal precipitates (molybdenum [Mo], ruthenium [Ru], palladium [Pd], technetium [Tc], and rhodium [Rh]) and the Xe fission gas atoms in UO2, by using density functional theory (DFT) in combination with the Hubbard $U$ correction to compute the formation energies involved. All DFT+$U$ calculations were performed with occupation matrix control to ensure antiferromagnetic ordering of UO2. The calculated formation and binding energies of the Xe and solid fission products in the NMP reveal that these metal precipitates form stable pair structures with Xe. Notably, the formation energy of Xe-metal pairs is lower than that of the isolated single defects in all instances, with Pd and Mo showing the most favourable binding energy, likely accounting for the observed pair structure formation.
@@ I Introduction
Recently, we performed advanced characterization of a spent UO2 fuel sample with an average burnup was approximately 40 MWdkg-1, from Belgium Reactor 3 (BR3) [27, 28]. Our investigation revealed that the fission product Xe forms intriguing pair structures with NMP. However, the fundamental mechanism behind these intriguing pair structures remains unknown. Although several studies have explored the stability of the NMP, the interactions between NMP and fission gases remain poorly understood. To address this knowledge gap, we employed first-principles calculations to investigate possible defect pair formations in UO2. Theoretical studies have previously utilized DFT simulations to understand defect energetics and diffusion of various fission products in actinide dioxides [29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. For example, Yun et al.[29] investigated the atomic diffusion mechanism of Xe in UO2 through vacancy-assisted diffusion, calculating the incorporation, binding, and migration energies using DFT with spin-orbit coupling (SOC). They found that an oxygen vacancy lowers the migration energy of a uranium vacancy and suggested that the strain energy of Xe significantly contributes to the clustering of vacancies, driving the vacancy-assisted diffusion of Xe in UO2. Nerikar et al.[30] used DFT+$U$ calculations to study the stability of neutral and charged intrinsic point defects, finding that predicted equilibrium properties and defect formation energies for neutral defect complexes align well with experimental trends, though values for charged complexes were lower than measured. Thompson et al. [33] conducted DFT+$U$ calculations to explore the energetics of various defects in UO2, including noble gases (He, Ne, Ar, Kr, Xe), Schottky defects, and the interaction between these defects. Andersson et al. [32] proposed mechanisms governing fission gas evolution by analyzing Xe solution thermodynamics, migration barriers, and the interaction of dissolved Xe atoms with uranium, demonstrating that Xe diffusion predominantly occurs via a vacancy-mediated mechanism. Hong et al. [34] used DFT to investigate the solubility and clustering of Ru fission products in UO2 suggesting that metallic dimers in Schottky defects likely form the nucleus of metallic precipitates. Cooper et al. [35] integrated DFT calculations of defect energies with empirical potential calculations of defect vibrational entropy to analyze the point defect concentrations as functions of oxygen partial pressure, temperature, and composition in UO2±x. Shilpa et al. [37] recently examined the formation and migration of various neutral defects such as vacancies, interstitials, antisites, Schotkky defects, and Frenkel defects in actinide oxides. Building on the work by Hong et al. [34] on the solubility and clustering of Ru fission products in UO2, we utilized the DFT+$U$ approach to understand the formation of various metal/fission-gas pair structures. We specifically characterized the formation energies, binding energy, and charge transitions of each Xe-metal pair structure, for metals of Tc, Mo, Pd, Ru, and Rh. Although it is well-established that charged defect play a crucial role in predicting specific experimental observations of UO2 under certain irradiation conditions [31, 30, 35, 38], this study focuses exclusively on neutral defects, because the focus of this work centers on formation energies and binding energies, and it is recognized that charged defects generally lowers the predicted energy values compared to neutral defects [30]. Consequently, while restricting our analysis to neutral defects may result in slight variations in the energy values, but will not comprise the overall conclusion of this study. For the 5MPs, our DFT+$U$ calculations identified the elements that are most energetically favorable to forming these pair structures. We anticipate that our results will inform future discussions regarding the impact of defect formation on the performance and stability of UO2. Furthermore, we expect that this work will stimulate future studies under various conditions encountered in nuclear reactors, ultimately advancing our understanding of defect dynamics in nuclear materials.
@@ II Experimental Details
Fig. 1(a) shows the low mag TEM image of pair structure between fission Xe/Kr gas bubbles (bright features) and NMP (M, dark features). Figs. 1(b) and (c) clearly illustrates the atomic-scale structure of a pair, with both Xe/Kr and NMP having a size of approximately 7 nm. The bright contrast observed in the bright-field image and the dark contrast in the high-angle annular dark-field image (i.e., Z-contrast) indicate a lower atomic density in the fission gas bubble as compared to the surrounding UO2 matrix. The embedded NMP disrupts the UO2 matrix, leading to lattice strain, which affects the contrast of the U atom columns (Figs. 1(b)). Though the crystal structure of the NMP is not fully resolved, it does not form a fully coherent interface with the UO2 matrix.
Atom Probe Tomography (APT) analysis was conducted on UO2 fuel to elucidate microstructural changes and the distribution of fission products. Detailed information on the APT instrument, sample preparation, and 3D reconstruction methodology is provided in the supplementary information (SI). The APT analysis site was strategically positioned adjacent to the TEM site to ensure that the same region of interest was utilized for both microstructural and chemical analysis.
Figure 2: (a) Representative 3D reconstruction of the UO2 tip, (b) distribution of Mo-Tc-Ru-Rh-Pd and an iso-concentration contour of Xe in a sub-volume, (c) proxigram from precipitates taken from figure (a)
Fig. 2(a) presents a representative 3D reconstruction image of a tip, featuring an iso-concentration map that illustrates Mo-Tc-Ru-Rh-Pd precipitates uniformly distributed throughout the analyzed volume, with sizes ranging between 2-3 nm in radius. Fig. 2(b) depicts a sub-volume of a 50 nm x 50 nm x 10 nm cube, highlighting the distribution of Mo-Tc-Ru-Rh-Pd and an iso-concentration contour of Xe. Fig. 2(c) displays a proximity histogram (proxigram) plot for all precipitates, facilitating the analysis of the chemistry across all precipitates shown in Figure 2(a). It is important to note that local magnification effects in APT may impact the quantitative analysis of the size, shape, and chemistry of the five metal precipitates and Xe, due to differences in the field evaporation behavior of the elements in the UO2 matrix [41, 42, 43]. Nonetheless, the concentrations of Mo, Pd, and Ru are the most significant. The APT analysis reveals distinct segregation of the five metal precipitates along with Xe, corroborating the findings from TEM analysis.
@@ III Simulation details
The chemical potential $\mu_{i}$ includes the zero-point vibrational energy and the temperature and pressure dependence of the chemical potential. By neglecting the zero-point vibration and the pressure effects, the term $\mu_{i}$ can be ignored. This approximation is reasonable, particularly for defect elements at zero temperature [61, 62, 63, 34, 64]. However,it has to be noted that the oxygen chemical potential is important when considering the hyperstoichiometric UO2 [35]. Thus, for defect elements, the third term in Equation (1) is just the DFT-calculated energy for each species. Furthermore, at equilibrium, the chemical potentials of U and O are related by the formation enthalpy of bulk UO2:
@@ IV Results and Discussion
To understand the formation energy of Xe-metal pairs, we first calculated the formation energy of single defects. The single defects in UO2 were constructed by replacing one U atom with either a Xe, Mo, Ru, Rh, Pd, or Tc atom. The relaxed-structure schematics for the single-atom Xe and Mo substitutions are shown in Fig. 3 (a), with all other single-defect relaxed structures being provided in the SI. The formation energy in an O-rich environment, as calculated using Equation 1 for all single defects considered in this work, is shown in Fig. 4. Among all the single defects analyzed, the Mo atom exhibits the lowest formation energy, indicating that the smallest energy penalty is incurred when substituting Mo at a U site. This is followed by Tc, Ru, Rh, Pd, and Xe. It is notable that the formation energy for single Mo and Tc atoms in UO2 is negative, indicating a solution in the lattice. However, given the experimental evidence of NMP formation, it is reasonable to deduce that the presence of other metals makes the formation of a separate phase thermodynamically more favorable. This suggests a complex interaction that warrants further computational investigation using larger length scale modeling. Following the analysis of single defects, we calculated the formation energies of dimer defects of the same species, with two atoms of Xe, Mo, Ru, Rh, Pd, or Tc being substituted at two neighboring U sites. The calculated formation energies of the dimers are listed in Fig. 4 (b) (denoted as “coupled”), in comparison with the formation energies measured when the two defects are spaced infinitely far apart (“isolated”). The total formation energy of the two isolated defects was estimated by summing together the formation energies of each individual defect.
@@
Figure 4 (b) clearly demonstrates that coupled dimers are energetically more favorable than dispersed atoms. Specifically, the XeU-XeU, RuU-RuU, PdU-PdU, and TaU-TaU pairs show a reduction in formation energy when in close proximity. In contrast, the changes in formation energy for RhU-RhU and MoU-MoU are marginal. These results indicate that metal atoms tend to form metallic clusters in UO2 fuel, rather than existing as dispersed atoms. Finally, we computed the formation energies of defect pairs involving Xe and various metals (Mo, Ru, Rh, Tc, and Pd) and compared these with the formation energies pertaining to when the metal and Xe atoms are isolated, as illustrated in Fig. 4 (c). Our results demonstrate that the formation energies of Xe-metal pairs are consistently lower than those of isolated defects, suggesting Xe-metal pair defects to be energetically favorable. Among the metals studied, the Xe-Mo pair exhibited the lowest formation energy, followed sequentially by Tc, Ru, Pd, and Rh. Furthermore, we calculated the binding energy of Xe-metal pairs, as illustrated in Fig. 4 (d). The results suggest that Xe-metal interactions are energetically favorable, with the highest binding energy observed for the Xe-Pd pair, followed by the Xe-Mo pair. Charge transfer analysis indicated that the charge transfer between Pd and Xe is negligible. Consequently, we attribute the substantial binding energy in the Xe-Pd system primarily to minimal mismatch between the strain field, as Pd as the closest atomic radii as that of Xe, among the five metal precipitates studied. This could potentially reduce the overall energy of the system, leading to a higher binding energy. In the case of the Xe-Mo pair, the binding energy arises from a combination of local charge redistribution and the change of strain field. In summary, the strong binding energies for the Xe-metal pairs can be attributed to a synergistic effect of charge transfer and change of strain field upon the formation of the Xe-metal pairs. This finding indicates that, of the metals investigated in this study, Pd and Mo are the most likely candidates to form stable pair structures with Xe.
@@
Figure 5: (a) Calculated formation energies of single defects at the U site, in the following order: Xe, Mo, Ru, Rh, Pd, and Tc. (b) Pair formation energies of dimers of the same species (blue), as compared to when these defects are isolated (orange). (c) Pair formation energies of dimers of the Xe-metal (blue), as compared to when these defects are isolated (orange).(d) Binding energy calculated by taking the energy difference for separated and coupled Xe-metal pairs
@@ V Conclusion
In conclusion, we employed DFT+$U$ calculations to unravel the fundamental mechanism underlying the formation of Xe-M pair structures in spent UO2 fuel samples from BR3. Our study calculated the formation energies and binding energies of five Xe-metal pairs, both in close proximity and in isolation. The results revealed that the formation energies of Xe-metal pairs are consistently lower than those of single defects, indicating enhanced stability for paired configurations. The binding energy of the Xe-M pairs follows the order of Xe-Pd $>$ Xe-Mo $>$ Xe-Tc $>$ Xe-Ru $>$ Xe-Rh, clearly highlighting that Pd and Mo are the most favorable metal for forming stable structures with Xe among the five metals considered. We attribute the stability of the metal-Xe pair to a synergistic effect of charge transfer and the change of strain field upon the formation of the Xe-metal pairs. Also, the APT analysis suggests that the concentration of Mo, Pd and Ru is relatively higher in the NMPs, indicating that these elements likely govern the fundamental mechanisms. This elucidation of the pair formation mechanism significantly advances our understanding of fission product behavior in nuclear fuels. Future energy-dispersive X-ray spectroscopy analyses of the BR3 samples are planned to further substantiate these findings through experimental validation. These insights enhance our comprehension of defect interactions in nuclear fuel matrices.